OpenStudy (anonymous):

Terrence normally uses seven-eighths teaspoon of dried basil to make his own salad dressing. He decides to change the recipe and use less dried basil. He decides that the flavor is perfect when using a total of one-half teaspoon. Write an addition equation that could be used to determine the amount of basil, b, Terrence excluded. Solve for b. Describe your steps.

OpenStudy (texaschic101):

1/2 + b = 7/8 -- subtract 1/2 from both sides b = 7/8 - 1/2 b = 7/8 - 4/8 b = 3/8 teaspoons

OpenStudy (texaschic101):

any questions ?

OpenStudy (anonymous):

No but thank you

OpenStudy (texaschic101):

anytime :)